express the complement of an angle represented by (y+10) degrees.
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

express the complement of an angle represented by (y+10) degrees.
-----------
Complementary angles add to 90 degrees.
C + (y+10) = 90

Answer by jim_thompson5910(35256)

The complement would be 90 - (y+10) = 90 - y - 10 = 80 - y degrees